# Analysis of Hierarchical AoII over unreliable channel: A Stochastic Hybrid System Approach

### v1

@@ III SHSs Analysis over Noisy Channel: The M/M/1/1 queue revisited
The AoI of M/M/1/1 queue has been thoroughly studied under different queueing disciplines. e.g., [5, 41, 1]. However, the queue has never been examined using a context-aware hierarchical age process. In this section, we provide a systematic way to obtain the closed-form results of the average hierarchical AoII in (2) using SHSs. The results would be later compared with the average AoI in Section V.
@@ III-A SHSs Formulations with Hierarchical AoII
To begin, we first need to characterize the system state, i.e., $\mathbf{q}\in\mathbb{Q}=\{0_{c},0_{e},1_{m},1_{u}\}$. State $0_{c}$ is defined as the correct (content-matched) state, where the AoII would not increase hence $\mathbf{\dot{x}}=[0,0]=\mathbf{0}_{n}$. State $1_{m}$ is defined as the mismatched state, where a content-changed update arrives and starts transmission. The AoII $\mathbf{x}(t)$ grows as the traditional AoI in this state, i.e., $\mathbf{\dot{x}}=[1,1]=\mathbf{1}_{n}$. State $0_{e}$ is defined as the error state, where no update is on transmission yet the content between ends is mismatched due to the channel noise. Hence, the age processes $\mathbf{x}(t)$ keep growing in $0_{e}$. State $1_{u}$ is the urgent state, where the transmitter is delivering a more important update. According to the model assumptions, $1_{u}$ can be reached when:
i). a newly sampled update with changed content preempts transmission. In that case, the content of $\hat{X}_{t}$ would lag at least two versions behind the transmitting update;
@@
The hierarchy of (2) can be implemented by adopting different age functions $f$ in each state, where the level of age dissatisfaction at states $0_{e}$ and $1_{u}$ is considered to be severer for the following two reasons.
@@
$\bullet$ The content of update being transmitting at $1_{u}$ is urged by the monitor due to conditions i) and ii) above, which also exacerbates the ageing of $\hat{X}_{t}$.
Consequently, we construct hierarchy schemes with four levels through the state-dependent derivative $f^{\prime}$ using two distinct piecewise functions.
- Linear hierarchy:
$f_{1}^{\prime}(\mathbf{q},\mathbf{x},t)=\begin{cases}\mathbf{0}_{n},\quad\textrm{if}\ \mathbf{q}=0_{c},\\ \mathbf{1}_{n},\quad\textrm{if}\ \mathbf{q}=1_{m},\\ m_{1}\mathbf{1}_{n},\quad\textrm{if}\ \mathbf{q}=0_{e},\\ m_{2}\mathbf{1}_{n},\quad\textrm{if}\ \mathbf{q}=1_{u}.\\ \end{cases}$ (5)
The parameters $m_{1}\geqslant 1$ and $m_{2}\geqslant 1$ are slopes representing two different levels of age dissatisfaction in the linear case.
- Quadratic hierarchy:
$f_{2}^{\prime}(\mathbf{q},\mathbf{x},t)=\begin{cases}\mathbf{0}_{n},\quad\textrm{if}\ \mathbf{q}=0_{c},\\ \mathbf{1}_{n},\quad\textrm{if}\ \mathbf{q}=1_{m},\\ k_{1}\mathbf{x},\quad\textrm{if}\ \mathbf{q}=0_{e},\\ k_{2}\mathbf{x},\quad\textrm{if}\ \mathbf{q}=1_{u}.\\ \end{cases}$ (6)
The parameters $k_{1}$ and $k_{2}$, where $k_{1}<k_{2}$, are the growth constants controlling the stability of SHSs, which would later be constrained by the state transition rates.
$\mathbf{q}=1_{u}$$\mathbf{\dot{x}}=k_{2}\mathbf{x}$$p\lambda$$p_{e}\mu$$\lambda$$\lambda$$p_{c}\mu$$\mathbf{q}=0_{e}$$\mathbf{\dot{x}}=k_{1}\mathbf{x}$$q\lambda$$p_{e}\mu$$\mathbf{q}=1_{m}$$\mathbf{\dot{x}}=\mathbf{1_{n}}$$q\lambda$$p\lambda$$p_{c}\mu$$\mathbf{q}=0_{c}$$\mathbf{\dot{x}}=\mathbf{0}_{n}$
Figure 3: Graphical representation of SHSs: the M/M/1/1 queue with preemption over noisy channel. The hierarchical growth of AoII takes the quadratic case for example, which is distinguished by the color of state. The mappings $\phi_{\mathit{l}}$ is omitted for simplicity.
@@
The subsequent results and stability analysis are more general using (5) or (6), which can be trivially simplified to a three-level case by letting $m_{1}=m_{2}$ or $k_{1}=k_{2}$. In addition, the hybrid of linear function at state $0_{e}$ and quadratic functions at state $1_{u}$ also follows basically the same analysis procedure as the quadratic case. Therefore, we take (6) as a common example.
To move forward, we now explain each transition $l$ with a constant intensity $\lambda_{\mathit{l}}(\mathbf{q},\mathbf{x})=\lambda_{l}$, which is illustrated in Fig. 3 with the pre-defined SHSs states.
$l=0,3,7$: Since the content of a newly sampled update remains unchanged with probability $q$ in the self-transition at state $0_{c}$, we assume that this kind of update would be maintained at the transmitter subject to an exponential random interval with rate $\mu$ (then dropped) instead of being transmitted. In the self-transition at state $1_{m}$, the preemption of an unchanged update also has no effect on the AoII. Moreover, since we consider the four-level hierarchy in the measure of timeliness for the noisy channel case, the state $1_{u}$ would stay under any preemption while the ageing process reaches the fourth level, which is the highest, at $1_{u}$.
$l=1,5$: An update with changed content leads to the second-level dissatisfaction between ends and the AoII starts to grow after the state is transferred into $1_{m}$ by $l=2$. Similarly, the update under transmission would also be preempted (then discarded) by a newly sampled update with changed content, which leads to the fourth-level dissatisfaction after the state is transferred into $1_{u}$ by $l=5$.
$l=2,6$: The content between ends is correctly matched after a successful transmission over the noisy channel. Meanwhile, the AoII is reset to zero by transitions $2,6$, respectively.
$l=4,8$: If the update is not successfully decoded by the monitor due to noise, the system state would be transferred into the error state $0_{e}$ by $l=4,8$. Then, $\hat{X}_{t}$ at the monitor suffers from a severer age dissatisfaction, namely, the third level in our consideration, due to the mismatch of content.
$l=9$: Any newly sampled update $X_{t}$ starts a transmission in $1_{u}$ when the system was in the error state before. Without loss of generality, we assume that the system goes through different out-of-sync situation between $1_{u}$ and $0_{s}$.
These state transitions are summarized in Table I, along with the possible changes of AoII after mappings $\{\phi_{\mathit{l}}:l\in\mathcal{L}\}$.
TABLE I: State Transitions and Changes of AoII over Noisy Channel
$1$ $0_{c}\to 1_{m}$ $p\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{c}}\end{bmatrix}$
$2$ $1_{m}\to 0_{c}$ $p_{c}\mu$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$3$ $1_{m}\to 1_{m}$ $q\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{m}}\end{bmatrix}$
$4$ $1_{m}\to 0_{e}$ $p_{e}\mu$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{m}}&v^{2}_{1_{m}}\end{bmatrix}$
$5$ $1_{m}\to 1_{u}$ $p\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{m}}\end{bmatrix}$
$6$ $1_{u}\to 0_{c}$ $p_{c}\mu$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$7$ $1_{u}\to 1_{u}$ $\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{u}}\end{bmatrix}$
$8$ $1_{u}\to 0_{e}$ $p_{e}\mu$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{u}}&v^{2}_{1_{u}}\end{bmatrix}$
$9$ $0_{e}\to 1_{u}$ $\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{e}}\end{bmatrix}$
III-B Generator for the SHSs
@@
as the expected value of AoII in each state $\overline{\mathbf{q}}$ at time $t$, where $\mathbf{v}_{\overline{\mathbf{q}}}(t)=[v^{1}_{\overline{\mathbf{q}}}(t),v^{2}_{\overline{\mathbf{q}}}(t)]$ tracks two different AoII process. Furthermore, we define the stationary solution (if exists) when $t\to\infty$ to be $\mathbf{v}_{\overline{\mathbf{q}}}=[v^{1}_{\overline{\mathbf{q}}},v^{2}_{\overline{\mathbf{q}}}]=\lim\nolimits_{t\to\infty}\mathbf{v}_{\overline{\mathbf{q}}}(t)$,
Given that $\psi$ is specified and $g(\mathbf{q},\mathbf{x},t)=0$, we introduce one more definition of an operator $L$ on the test function $\psi$[42], that is,
@@
which is called the extended generator for our SHSs. $f^{\prime}(\mathbf{q},\mathbf{x},t)$ is in the form of (5) or (6), and $\lambda_{l}$ for the noisy channel case is given in Table I. Note that the extended generator $(L\psi)_{0_{c}}=(L\psi)_{0_{c}}(\mathbf{q},\mathbf{x},t)$ at state $0_{c}$ is zero since the AoII at $0_{c}$ never grows at all.333 As noted in [9], $0_{c}$ is called the irrelevant state to AoII. However, this kind of states does affect the stationary distributions.
@@
We now can leverage (10) to derive the closed-form expressions of average AoII for the M/M/1/1 queue under linear hierarchy $f_{1}^{\prime}(\mathbf{q})$, which is concluded as [1, Theorem 1]. However, in order to calculate the average AoII under quadratic hierarchy $f_{2}^{\prime}(\mathbf{q})$, we need to modify the aforementioned theorem. Hence, we summarize our first result in the following.
@@
(i). The finite-state SHSs Markov Chain is ergodic with stationary distributions $\{\pi_{\overline{\mathbf{q}}}:\overline{\mathbf{q}}\in\mathbb{Q}\}$ given by the global balance equations:
@@
(ii). For linear growth, the expected value $\mathbf{v}_{\overline{\mathbf{q}}}$ can be calculated by
@@
where the age grows as a linear function at $\overline{\mathbf{q}}$, and $\mathbf{q}_{l}$ is denoted as the other state in $\mathcal{L}^{i}_{\overline{\mathbf{q}}}$.
(iii). For quadratic growth, the expected value $\mathbf{v}_{\overline{\mathbf{q}}}$ can be calculated by
@@
where the age grows as a quadratic function at $\overline{\mathbf{q}}$, and $k_{\overline{\mathbf{q}}}$ is the growth constant associated with $\overline{\mathbf{q}}$.
@@
$x=\sum_{\overline{\mathbf{q}}\in\mathbb{Q}}\mathbf{v}_{\overline{\mathbf{q}}}^{2},$ (15)
for different functions $f^{\prime}(\overline{\mathbf{q}})$.
The matrix $\mathbf{A}_{l}$ in (13) and (14) is the transition matrix associated with the mapping $\phi_{\mathit{l}}$, which is also given in Table I. Since the first two parts of Theorem 1 have been proved in [1], we omit here for simplicity. The proof of (iii) appears in Appendix A. Note that (14) is only valid under some constrained parameters $k_{1}$ and $k_{2}$, which is called the stability conditions for quadratic hierarchy and is discussed later in Section III-D.
@@
Theorem 1 gives the general procedure to calculate the average AoII for different hierarchy schemes. First, we need to calculate the stationary distribution for each finite-state SHSs Markov chains using (12), the existence of which ensures the ergodicity. Then, the average AoII can be calculated separately at each state $\mathbf{q}$ using different state-dependent growth function $f^{\prime}(\mathbf{q})$, as derived in (13) and (14). Hence, our theorem can be easily generalized to multi-level hierarchy of state-dependent AoII using any hybrid of linear and quadratic functions. Meanwhile, the theorem can also be further extended to derive the stationary moments (also MGF) of AoII as did in [1, Theorem 1] under the same stability conditions.
III-C Average AoII under Linear Hierarchy

### v2

@@ III SHSs Analysis over Noisy Channel: The M/M/1/1 queue revisited
The AoI of the M/M/1/1 queue has been thoroughly studied under different queueing disciplines. e.g., [2, 41, 1]. However, the queue has never been examined using a context-aware hierarchical AoII (ageing) process. In this section, we provide a systematic way to obtain the closed-form results of the average hierarchical AoII in (4) using SHSs. The results would be later compared with the average AoI in Section V.
@@ III-A SHSs Formulations with Hierarchical AoII
To begin, we first need to characterize the system state, i.e., $\mathbf{q}\in\mathbb{Q}=\{0_{c},0_{e},1_{m},1_{u}\}$, under the noisy channel case. State $0_{c}$ is defined as the correct (content-matched) state, where the AoII would not increase hence $\mathbf{\dot{x}}=[0,0]=\mathbf{0}_{n}$. State $1_{m}$ is defined as the mismatched state, where a content-changed update arrives and starts transmission. The AoII $\mathbf{x}(t)$ grows as the traditional AoI in this state, i.e., $\mathbf{\dot{x}}=[1,1]=\mathbf{1}_{n}$. State $0_{e}$ is defined as the error state, where no update is on transmission yet the content between ends is mismatched due to the channel noise. Hence, the age processes $\mathbf{x}(t)$ keep growing in $0_{e}$. State $1_{u}$ is the urgent state, where the transmitter is delivering a more important update. According to the model assumptions, $1_{u}$ can be reached when:
i). A newly sampled update with changed content preempts transmission. In that case, the content of $\hat{X}_{t}$ lags at least two versions behind the transmitting update;
@@
Compared to states $0_{c}$ and $1_{m}$, the level of age dissatisfaction at states $0_{e}$ and $1_{u}$ is considered to be severe for the following two reasons.
@@
$\bullet$ The content of update being transmitting at $1_{u}$ is urged by the monitor due to issues i) and ii) given above, which also exacerbates the ageing of $\hat{X}_{t}$.
Based on these considerations, the AoII processes at the monitor should not be the same under different system states, in other words, hierarchical AoII is needed. To that end, we construct two different hierarchy schemes, each with four hierarchical levels, using the growth rate function $f^{\prime}$.
- Hierarchy Scheme 1:
[…]
Scheme 1 represents the case that $f^{\prime}$ has four levels of constant growth rates at each state. The parameters $m_{2}\geqslant m_{1}\geqslant 1$ are slopes representing two different levels of age dissatisfaction in this case.
- Hierarchy Scheme 2:
[…]
Scheme 2 represents the case that $f^{\prime}$ has two levels of constant growth rates and two levels of linear growth rates in $\mathbf{x}$. The parameters $k_{1}$ and $k_{2}$, where $k_{2}\geqslant k_{1}$, are the growth constants controlling the stability of SHSs, which will later be used in obtaining the stability conditions of AoII SHSs.
@@
The subsequent results using (7) or (8) can be trivially simplified to a three-level hierarchical case by letting $m_{1}=m_{2}$ or $k_{1}=k_{2}$. In addition, if we change the level of age dissatisfaction at state $1_{u}$ in scheme 2 into the level of constant growth rate, the analysis of such AoII SHSs is basically the same as scheme 2. Therefore, we take (8) as a common example in this work.
III-B Generator, Positive Recurrence and Calculation of SHSs
@@
as the expected value of AoII in each state $\overline{\mathbf{q}}$ at time $t$, where $\mathbf{v}_{\overline{\mathbf{q}}}(t)=[v^{1}_{\overline{\mathbf{q}}}(t),v^{2}_{\overline{\mathbf{q}}}(t)]$ tracks two different AoII process. Furthermore, we define the stationary solution (if exists) when $t\to\infty$ to be $\mathbf{v}_{\overline{\mathbf{q}}}=[v^{1}_{\overline{\mathbf{q}}},v^{2}_{\overline{\mathbf{q}}}]=\lim\nolimits_{t\to\infty}\mathbf{v}_{\overline{\mathbf{q}}}(t)$.
Given that $\psi$ is specified and $g(\mathbf{q},\mathbf{x},t)=0$, we introduce an operator $L$ on the test function $\psi$[42], that is,
@@
which is called the extended generator for our SHSs. $f^{\prime}(\mathbf{q},\mathbf{x},t)$ is in the form of (7) or (8), and $\lambda_{l}$ for the noisy channel case is given in Table I. Note that the extended generator $(L\psi)_{0_{c}}=(L\psi)_{0_{c}}(\mathbf{q},\mathbf{x},t)$ at state $0_{c}$ is zero since the AoII at $0_{c}$ never grows at all. [^4] As noted in [6], $0_{c}$ is called the irrelevant state to AoII. However, this kind of states does affect the stationary distributions of discrete state $\mathbf{q}$.
@@
Positive Recurrence
To move forward, we need to establish the weak-sense stability, i.e., positive recurrence, for our AoII SHSs under both hierarchy schemes.
Lemma 2.
The AoII SHSs under the noisy channel case are positive recurrent in both hierarchy schemes.
The proof of Lemma 2 appears in Appendix B, which can also be used in proving the positive recurrence of AoII SHSs under the collision channel case. Lemma 2 also guarantees that the AoII SHSs are regular under both hierarchy schemes where the expected AoII $\mathrm{E}[\mathbf{x}(t)]$ is always finite with probability one. Moreover, we can establish the ergodicity of the AoII SHSs based on the positive recurrence using the same technique introduced in [35] to prove steady-state convergence of average AoII [^5] The technique of proof has been revealed in our work [43, Theorem 2], hence it is omitted here for simplicity.. With these, we can now proceed with our calculation for closed-form expressions of average AoII for the M/M/1/1 queue under $f_{1}^{\prime}(\mathbf{q})$, which basically follows [1, Theorem 1]. However, in order to calculate the average AoII under $f_{2}^{\prime}(\mathbf{q})$, we need to modify the aforementioned theorem. Hence, we summarize our first result in the following.
@@
(i). The stationary distributions $\{\pi_{\overline{\mathbf{q}}}:\overline{\mathbf{q}}\in\mathbb{Q}\}$ of discrete states in SHSs are given by the global balance equations:
@@
(ii). If the growth rate function $f^{\prime}$ is constant at state $\overline{\mathbf{q}}$, the expected value $\mathbf{v}_{\overline{\mathbf{q}}}$ can be calculated by
@@
where $\mathbf{q}_{l}$ is denoted as the other state in $\mathcal{L}^{i}_{\overline{\mathbf{q}}}$.
(iii). If the growth rate function $f^{\prime}$ is linear in $\mathbf{x}$ at state $\overline{\mathbf{q}}$, the expected value $\mathbf{v}_{\overline{\mathbf{q}}}$ can be calculated by
@@
where $k_{\overline{\mathbf{q}}}$ is the growth constant associated with $\overline{\mathbf{q}}$.
@@
$x=\sum_{\overline{\mathbf{q}}\in\mathbb{Q}}v_{\overline{\mathbf{q}}}^{2}.$ (17)
The matrix $\mathbf{A}_{l}$ in (15) and (16) is the transition matrix associated with the linear mapping $\phi_{\mathit{l}}$. Since the first two parts of Theorem 1 have been proved in [1], we omit here for simplicity. The proof of (iii) appears in Appendix A. Note that (16) is only valid under some constrained parameters $k_{\mathbf{q}}$, which, in this work, is called the steady-state stability conditions of the AoII SHSs and is discussed in Section III-D.
@@
Theorem 1 gives the general procedure to calculate the average AoII for different hierarchy schemes. First, we need to calculate the stationary distribution for each finite-state SHSs Markov chains using (14), the existence of which ensures the ergodicity. Then, the average AoII can be calculated separately at each state $\mathbf{q}$ using different state-dependent growth function $f^{\prime}(\mathbf{q})$, as derived in (15) and (16). Hence, our theorem can be easily generalized to multi-level hierarchy of state-dependent AoII using any hybrid of constant and linear (in $\mathbf{x}$) growth rate functions.
III-C Average AoII under Hierarchy Scheme 1
[…]
Figure 3: Graphical representation of SHSs: the M/M/1/1 queue with preemption over noisy channel. The hierarchical growth of AoII takes the second case as an example, which is distinguished by the color of state.
We now explain each transition $l$ with a constant $\lambda_{\mathit{l}}(\mathbf{q},\mathbf{x})=\lambda_{l}$ under the noisy channel, which is illustrated in Fig. 3 and summarized in Table I with the pre-defined SHSs states.
TABLE I: State Transitions and Changes of AoII over Noisy Channel
11 $0_{c}\to 1_{m}$ $p\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{c}}\end{bmatrix}$
22 $1_{m}\to 0_{c}$ $p_{c}\mu$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
33 $1_{m}\to 1_{m}$ $q\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{m}}\end{bmatrix}$
44 $1_{m}\to 0_{e}$ $p_{e}\mu$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{m}}&v^{2}_{1_{m}}\end{bmatrix}$
55 $1_{m}\to 1_{u}$ $p\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{m}}\end{bmatrix}$
66 $1_{u}\to 0_{c}$ $p_{c}\mu$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
77 $1_{u}\to 1_{u}$ $\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{u}}\end{bmatrix}$
88 $1_{u}\to 0_{e}$ $p_{e}\mu$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{u}}&v^{2}_{1_{u}}\end{bmatrix}$
99 $0_{e}\to 1_{u}$ $\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{e}}\end{bmatrix}$
$l=0,3,7$: Since the content of a newly sampled update remains unchanged with probability $q$ in the self-transition at state $0_{c}$ ($l=0$), we assume that this kind of update would be maintained at the transmitter subject to an exponential random interval with rate $\mu$ (then dropped) instead of being transmitted. In the self-transition at state $1_{m}$ ($l=3$), the preemption of an unchanged update also has no effect on the AoII. Moreover, since we consider the four-level hierarchy in the measure of timeliness for the noisy channel case, the state $1_{u}$ would stay under any preemption while the AoII process reaches the fourth level, which is the highest, at $1_{u}$ ($l=7$).
$l=1,5$: An update with changed content leads to the second-level dissatisfaction between ends and the AoII starts to grow after the state is transferred into $1_{m}$ ($l=1$). Similarly, the update under transmission would also be preempted (then discarded) by a newly sampled update with changed content, which leads to the fourth-level dissatisfaction after the state is transferred into $1_{u}$ ($l=5$).
$l=2,6$: The content between ends is matched after a successful transmission over the noisy channel. Meanwhile, the AoII is reset to zero by these two transitions, respectively.
$l=4,8$: If the update is not successfully decoded by the monitor due to noise, the system state would be transferred into the error state $0_{e}$. Then, the system suffers a severe age dissatisfaction, namely, the third-level in our consideration, due to the mismatch of content.
$l=9$: If the system were in the error state $0_{e}$, any newly sampled update $X_{t}$ transfers the system state into $1_{u}$ and starts a transmission.